ABC Company is considering whether a project requiring the purchase of new equipment worth investing. The firm spent $20,000 three months ago to conduct market study. The cost of a new machine is $160,000, and the firm has to spend additional $10,000 to get it shipped and installed. This project will increase annual revenues by $225,000 and annual costs by $45,000. If the firm undertakes this project, $50,000 in net working capital investment is required. What is the initial outlay of this project? Round to the nearest penny. Do not include a dollar sign in your answer. Type the absolute value of the answer. (i.e. If your answer is -$20,000 since it is cost to replace, type 20000 as your answer.)
Answer #1

Initial outlay of the project = Initial investment + Investment in working capital

Initial investment = cost of machine + shipping costs

= 160000 + 10000 = 170000

Initial outlay = 170000 + 50000

= 220000